5+(c/9) =(-31)
c/9 is a fraction and I cant seem to figure out what number c=????? HELP!

Answer by checkley77(12844)

5+C/9=-31
C/9=-31-5
C/9=-36
C=-36*9
C=-324 ANSWER.
PROOF:
5+(-324/9)=-31
-324/9=-31-5
-36=-36